South Company sells a single product for $20 per unit. If variable expenses are 60% of sales and fixed expenses total $9,600, the break-even point will be:
Answer #1

Contribution margin=Sales-Variable cost  

=(100-60)=40%

Breakeven=Fixed expenses/Contribution margin ratio

=(9600/0.4)

=$24000

=(24000/20)

=1200 units